OpenStudy (anonymous):

11. Apply the division property of equality to find the next step in solving this equation: 3x/4=12

OpenStudy (anonymous):

3x = 4*12

OpenStudy (texaschic101):

are you sure its not asking for the multiplication property of equality ?

OpenStudy (anonymous):

I'm not sure. That's all is says in the quetion

OpenStudy (texaschic101):

this is how I would solve it.. 3x/4 = 12 --- multiply both sides by 4 (multiplication property) 3x = 12 * 4 3x = 48 -- divide both sides by 3 (division property) x = 48/3 x = 16 check.. 3(16) / 4 = 12 48/4 = 12 12 = 12 (correct)

OpenStudy (anonymous):

Oh I see what you mean